# SpeechRole: A Large-Scale Dataset and Benchmark for Evaluating Speech Role-Playing Agents

### v4

@@ Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
[…]
@@ 1 Introduction
Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [DBLP:conf/iui/ChangSZJ0K23]. Although preliminary studies have explored role-playing incorporating visual or speech modalities [DBLP:conf/iclr/DaiH0J0025, DBLP:journals/corr/abs-2505-20277], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs). SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions. In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
To address this gap, we propose SpeechRole, a unified framework for the development and evaluation of SRPAs (see Figure 1). We first introduce SpeechRole-Data, a large-scale dataset comprising 112k speech-based dialogues covering 98 diverse roles. Each role features distinctive timbre and prosodic characteristics, enabling expressive and role-consistent speech interactions.
Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses agents across three dimensions: interaction competence, speech expressiveness, and role consistency. This benchmark supports both cascaded and end-to-end architectures, facilitating comprehensive and comparable analyses across different paradigms.
Our experiments reveal complementary strengths and weaknesses between SRPA paradigms. Cascaded systems offer strong modular control and stable role identity across interactions, but may suffer from cross-module mismatches or limited expressiveness. In contrast, end-to-end models produce more coherent and natural speech, yet often struggle with speaker consistency and preserving character traits over extended dialogues. These trade-offs highlight the unique challenges in building effective SRPAs that balance interaction ability, speech expressiveness, and role-playing fidelity.
Overall, our main contributions are as follows:
@@
    We construct SpeechRole-Data, a large-scale multilingual corpus covering 98 roles and 112k speech-based dialogues with rich prosodic and timbral variation.
@@
    We develop SpeechRole-Eval, a comprehensive benchmark designed to evaluate SRPA capabilities across key dimensions including interaction ability, speech expressiveness, and role-playing fidelity.
@@
    We systematically compare cascaded and end-to-end paradigms, revealing their strengths, limitations, and offering insights for future SRPA development.
@@ 2 Related Work
Role-Playing Agents Across Modalities.
Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality. ChatHaruhi [DBLP:journals/corr/abs-2308-09597] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], Character-LLM [DBLP:conf/emnlp/ShaoLDQ23], and RoleFact [DBLP:conf/emnlp/SadeqXKLGM24] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations. Beyond text, MMRole [DBLP:conf/iclr/DaiH0J0025] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the value of speech for real-time, voice-aware character interaction.
To assess role-playing capabilities, benchmarks such as RoleEval [DBLP:journals/corr/abs-2312-16132], RoleInteract [DBLP:journals/corr/abs-2403-13679], SocialBench [DBLP:conf/acl/ChenCYXXSQLZH24], and CharacterEval [DBLP:conf/acl/TuFTSSGY24] evaluated behavioral consistency and persona fidelity, while INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24] and ECHO [DBLP:journals/corr/abs-2404-13957] introduced human-aligned, psychometric evaluation paradigms. These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity. However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
Interactive Speech Agents.
In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24]. Modern TTS models like F5TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], E2-TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], Spark-TTS [DBLP:journals/corr/abs-2503-01710], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025], and XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24] have significantly improved output quality, expressiveness, and controllability. However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
End-to-end systems aim to unify speech perception and generation within a single model. Early efforts like SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [DBLP:journals/corr/abs-2306-12925] utilized vocoder-based representations for continuous speech modeling. Recent advances, including Moshi [DBLP:journals/corr/abs-2410-00037], LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025], and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
Figure 2: Examples of the three major task categories in SpeechRole.
@@ 3 SpeechRole-Data
We introduce SpeechRole-Data, a large-scale dataset for speech-based role-play, featuring 98 roles and over 112k dialogues with reference audio and rich character profiles. To ensure personality and voice consistency, we follow a structured pipeline spanning role selection, dialogue generation, and audio processing. Details are in the appendix.
3.1 Text Data Construction
Role Selection.
Role Metadata Extraction.
Dialogue Generation.
Dialogue Filtering and Deduplication.
The generated data undergoes a multi-stage filtering process. First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese. To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
@@ 3.1 Text Data Construction
3.2 Role Voice Collection
Voice Collection and Preprocessing.
To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works. All audio is anonymized and fragmented into brief utterances to comply with fair use. Raw streams are extracted using ffmpeg [ffmpeg] and converted to mono 16 kHz WAV format.
Voice Cleaning and Annotation.
We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, including source separation, speaker diarization, and voice activity detection. This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23] and scored with DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22].
Figure 3: Speech duration distribution of users and roles.
Speaker Identification.
The diarization process outputs speaker indices without explicit identity information. To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to analyze transcribed text. The models’ integrated results are then manually verified.
Reference Audio Ranking and Selection.
To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity. We use the CAM++ model [DBLP:conf/interspeech/WangZCC023] to extract a speaker embedding for each clip. The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
@@ 3.3 Dataset Statistics
Dataset Overview.
SpeechRole-Data comprises 98 distinct roles and 112k dialogue samples, including both single-turn and multi-turn conversations (see Table 1). Each role is accompanied by a detailed role profile and a reference audio clip that demonstrates its unique speaking style, such as timbre, rhythm, and prosody. Each dialogue sample contains both text and speech, along with contextual background information relevant to the role, thereby supporting realistic, evidence-based role-playing. During the evaluation phase, the test set consists of two single-turn and two multi-turn dialogues sampled for each role. The remaining data are divided into a training set (78 roles) and a validation set (20 roles). Detailed examples are provided in Figure 2.
Speech Duration.
We analyze the speech duration of users and roles. As shown in Figure 3, the duration of role utterances is generally longer, typically ranging from 10 to 20 seconds. This is mainly due to the narrative or expressive nature of their content. In contrast, user utterances are concentrated within 3 to 6 seconds, reflecting their concise and goal-oriented characteristics.
@@ 4 SpeechRole-Eval
To systematically assess SRPAs, we introduce SpeechRole-Eval, a comprehensive benchmark that evaluates model performance across three key dimensions: fundamental interaction ability, speech expressiveness, and role-playing fidelity. In contrast to prior work that primarily focuses on text, SpeechRole-Eval places special emphasis on the speech modality as a critical component of evaluation. The scoring process is fully automated, powered by a strong multimodal model, and has been validated against human judgments to ensure reliability. This section describes the task setup, evaluation metrics, and the overall evaluation pipeline.
4.1 Task Types
To assess the cognitive, affective, and social competencies of SRPAs, user prompts are categorized into three major task categories based on communicative intent and interactional focus: (1) Internal Reasoning, which concerns the character’s rich subjective world, including personal perspectives, emotional responses, and underlying motivations; (2) Experiential Narration, which emphasizes the character’s ability to vividly recall past events, describe concrete actions, and provide relevant background or historical information; and (3) Social Communication, which highlights interpersonal understanding achieved through nuanced evaluative judgments, the provision of advice, and relational reasoning in context. These categories reflect the core dimensions of role-playing dialogue and enable a comprehensive, structured assessment of role-playing abilities (see Figure 2).

### v6

@@ Abstract
Speech is essential for realistic role-playing, yet existing work on role-playing agents largely centers on text, leaving Speech Role-Playing Agents (SRPAs) underexplored and lacking systematic evaluation. We introduce SpeechRole, a unified framework for developing and assessing SRPAs. SpeechRole-Data contains 98 roles and 111k speech-to-speech conversations with rich timbre and prosodic variation, providing large-scale resources for training SRPAs. SpeechRole-Eval offers a multidimensional benchmark that directly evaluates generated speech, preserving paralinguistic cues and measuring interaction ability, speech expressiveness, and role-playing fidelity. Experiments show that end-to-end SRPAs such as GPT-4o Audio achieve strong fluency and naturalness, but remain limited in prosody consistency and emotion appropriateness. In contrast, current open-source end-to-end models exhibit substantial performance gaps across multiple evaluation dimensions. Cascaded and end-to-end systems achieve comparable results in interaction ability and role-playing fidelity, suggesting that these aspects are still largely influenced by the underlying text-based language models. We release all data, code, and evaluation tools at https://github.com/yuhui1038/SpeechRole.
*Equal Contribution.
Corresponding Author.
@@ 1 Introduction
Speech Role-Playing Agents (SRPAs) are speech-to-speech systems designed to generate spoken responses in character, reproducing distinctive vocal timbre, prosody, and persona traits while maintaining coherence across interactions. Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [DBLP:conf/iui/ChangSZJ0K23]. Although preliminary studies have explored speech-based role-playing [DBLP:journals/corr/abs-2505-20277], large-scale datasets and systematic evaluation benchmarks for SRPAs remain scarce.
To address this gap, we introduce SpeechRole, a unified framework for developing and evaluating SRPAs. SpeechRole-Data is a large-scale speech-to-speech dataset containing 111k dialogues across 98 roles, offering diverse vocal traits and prosodic patterns that support the construction of SRPAs. Building on this foundation, SpeechRole-Eval provides a multidimensional benchmark that directly evaluates generated speech without relying on speech-to-text conversion, thus preserving critical paralinguistic information. The benchmark assesses SRPAs along three complementary dimensions: interaction ability, speech expressiveness, and role-playing fidelity. Together, SpeechRole-Data and SpeechRole-Eval establish a foundation for systematic research on speech-based role-playing.
Our empirical findings highlight clear differences between cascaded and end-to-end SRPAs. The most advanced end-to-end system, GPT-4o Audio, demonstrates notable advantages in speech fluency and naturalness, while still exhibiting limitations in prosody consistency and emotion appropriateness. In contrast, current open-source end-to-end models show substantial performance gaps across multiple evaluation dimensions. Meanwhile, for interaction ability and role-playing fidelity, cascaded and end-to-end systems achieve comparable results, suggesting that these capabilities remain strongly influenced by the underlying text-based large language models rather than the specific speech modeling pipeline. Overall, our main contributions are as follows:
@@
    We construct SpeechRole-Data, a large-scale speech-to-speech corpus with 111k dialogues across 98 roles, offering diverse timbral and prosodic patterns for developing SRPAs.
@@
    We introduce SpeechRole-Eval, a multidimensional benchmark that directly evaluates generated speech and assesses interaction ability, speech expressiveness, and role-playing fidelity without relying on speech-to-text conversion.
@@
    We provide a systematic comparison of cascaded and end-to-end SRPAs, characterizing their strengths and limitations and analyzing how system design impacts speech quality and role-playing performance.
Figure 2: Examples of the three major task categories in SpeechRole.
@@ 2 Related Work
2.0.0.1 Role-Playing Agents Across Modalities.
Recent research has increasingly explored the role-playing capabilities of LLMs, with most efforts focusing on the text modality. Notable examples include ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], which propose methods to improve persona simulation and reduce hallucinations. OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the importance of incorporating speech for real-time, voice-aware character interactions. Existing benchmarks, such as RoleEval [DBLP:journals/corr/abs-2312-16132] and INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24], focus on behavioral consistency and persona fidelity. However, systematic evaluation frameworks for speech-based role-playing remain scarce, leaving a gap in assessing vocal expression, emotional depth, and user engagement.
2.0.0.2 Interactive Speech Agents.
Speech-interactive LLMs typically follow two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modular design has been widely adopted in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24]. Recent TTS models have improved naturalness, expressiveness, and controllability, although cascaded pipelines can suffer from error accumulation and limited cross-modal context integration. End-to-end systems aim to unify speech perception and generation within a single model. Early work such as SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies. More recent architectures, including LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025] and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], advance end-to-end speech modeling through dual-modality training, streaming capabilities, and low-latency generation, enabling more natural and efficient speech-driven interactions. Despite these advances, systematic evaluation of speech role-playing capabilities in end-to-end settings remains limited.
@@ 3 SpeechRole-Data
SpeechRole-Data is constructed to support SRPAs, which require speech-based dialogues that capture not only a character’s linguistic behavior but also their distinctive vocal attributes. Building such data necessitates two essential components: persona-consistent dialogue content that reflects how the character would interact, and representative voice samples that preserve their timbre, prosody, and speaking style.
To obtain both elements, we curate 98 characters drawn from television dramas, films, animations, and games. For each character, we collect narrative profiles and scripts. Based on these materials, an LLM is prompted to generate user–character dialogue texts that remain faithful to the original personality and storyline. In parallel, we extract clean speech segments from the corresponding audiovisual sources to capture each character’s timbre and prosody. Using the generated texts and collected voice samples, we synthesize full speech-to-speech dialogues with a TTS system. This pipeline produces a corpus of 111k spoken interactions tailored for SRPA development and evaluation.
@@
3.1 Text Data Construction
3.1.0.1 Role Selection.
3.1.0.2 Role Metadata Extraction.
3.1.0.3 Dialogue Generation.
3.1.0.4 Dialogue Deduplication.
To improve data diversity and avoid repeated conversational patterns, we perform dialogue-level deduplication using semantic similarity. For each dialogue, we compute its embedding-based similarity with all other dialogues using all-MiniLM-L6-v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese. Dialogues with similarity above 0.9 for English or 0.85 for Chinese are considered near-duplicates, and only one instance is retained.
3.2 Role Voice Collection and Synthesis
3.2.0.1 Voice Collection and Preprocessing.
To acquire authentic reference voices, we collect audio from sources such as the game Genshin Impact, whose assets permit non-commercial academic research use. For other audiovisual works, we only extract short speech fragments for analysis and dataset construction and do not distribute or reproduce any original copyrighted material. All collected audio is anonymized and segmented into brief utterances. Raw streams are extracted using ffmpeg [ffmpeg] and converted to mono 16 kHz WAV format.
3.2.0.2 Audio Cleaning and Segmentation.
We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, performing source separation, speaker diarization, and voice activity detection. This pipeline produces clean 3–10 second single-speaker clips. The quality of each clip is assessed using DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22], and clips with an overall quality score below 3 are filtered out.
3.2.0.3 Role-Level Speaker Identification.
Speaker diarization groups the original audiovisual audio into segments corresponding to different speakers, assigning each speaker a numeric ID but not revealing which character the ID represents. To map each segment to the correct character, we first transcribe the clips using whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23]. We then use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to infer character identity based on the transcribed text and contextual cues. Finally, all assignments are manually verified for accuracy.
3.2.0.4 Reference Voice Selection.
We select a representative reference voice for each role by ranking all its cleaned clips according to vocal consistency. For each clip, we extract a speaker embedding using the CAM++ model [DBLP:conf/interspeech/WangZCC023]. We then compute its average cosine similarity to all other clips belonging to the same role and choose the clip with the highest similarity as the reference voice, as it best reflects the role’s characteristic timbre and speaking style.
3.2.0.5 Role and User Speech Synthesis.
Using the collected reference voices, we synthesize role utterances with CosyVoice2 [DBLP:journals/corr/abs-2412-10117], which conditions on each role’s selected clip to preserve timbre and prosody. User utterances are generated with the VolcEngine TTS system [volcengine] using a single fixed voice, as the user does not require persona-specific vocal variation. This yields complete speech-to-speech dialogues aligned with both the textual content and role-specific vocal characteristics.
Figure 3: Speech duration distribution of SpeechRole-Data.
@@ 3.3 Dataset Statistics
3.3.0.1 Dataset Overview.
SpeechRole-Data comprises 98 distinct roles and 111k speech-to-speech dialogue samples, covering both single-turn and multi-turn conversations. Each role is paired with a detailed role profile and a reference audio clip that illustrates its characteristic speaking style, including timbre, rhythm, and prosody. Each dialogue sample contains speech data and contextual information relevant to the role, enabling realistic and evidence-grounded role-playing.
The dataset is divided into three subsets. The training set contains dialogues from 78 roles. The out-of-domain set includes 20 roles that do not appear in the training data, enabling evaluation of generalization to unseen characters. The test set covers all 98 roles, with four carefully curated dialogue samples per role, resulting in a total of 392 test instances. These instances are independently selected from both the training and out-of-domain pools (see Table 1). The scale of this test set is comparable to prior role-playing evaluation benchmarks, such as the 400-sample evaluation in DBLP:journals/corr/abs-2505-20277 and the 294-sample benchmark introduced by DBLP:conf/iclr/DaiH0J0025. Detailed examples are provided in Figure 2.
3.3.0.2 Speech Duration.
We analyze the duration of individual speech segments for both users and roles. As shown in Figure 3, single utterances from roles are generally longer, typically ranging from 10 to 20 seconds, reflecting their narrative or expressive nature. In contrast, user utterances are concentrated between 3 and 6 seconds, highlighting their concise and goal-oriented characteristics.
Models IA $C$ SF SN PC EA PeC KC Overall
English Evaluation Results
Qwen3-8B 0.926 0.955 0.861 0.825 0.845 0.884 0.871 0.908 0.884
Llama-3.1-8B 0.915 0.932 0.846 0.823 0.824 0.864 0.826 0.886 0.865
Mistral-7B 0.843 0.894 0.859 0.820 0.781 0.806 0.752 0.819 0.822
Alibaba Cloud API 0.895 0.954 0.823 0.786 0.768 0.799 0.784 0.867 0.835
LLaMA-Omni 0.791 0.865 0.880 0.793 0.674 0.668 0.659 0.734 0.758
Qwen2.5-Omni 0.537 0.647 0.682 0.640 0.464 0.490 0.403 0.595 0.557
GPT-4o Audio 0.901 0.990 0.942 0.912 0.811 0.846 0.807 0.881 0.886
Chinese Evaluation Results
Qwen3-8B 0.823 0.871 0.858 0.829 0.874 0.875 0.844 0.858 0.854
Llama-3.1-8B 0.818 0.776 0.820 0.822 0.785 0.798 0.714 0.748 0.785
Mistral-7B 0.717 0.700 0.840 0.811 0.712 0.702 0.602 0.664 0.719
Alibaba Cloud API 0.905 0.929 0.919 0.913 0.861 0.865 0.828 0.827 0.881
Qwen2.5-Omni 0.644 0.767 0.890 0.799 0.608 0.620 0.511 0.652 0.686
GPT-4o Audio 0.858 0.961 0.902 0.896 0.828 0.841 0.800 0.868 0.869
Table 2: Comprehensive evaluation results of 7 SRPAs on the SpeechRole-Eval benchmark. The table reports model performance across eight evaluation metrics: IA (Instruction Adherence), $C$ (Conversational Coherence), SF (Speech Fluency), SN (Speech Naturalness), PC (Prosody Consistency), EA (Emotion Appropriateness), PeC (Personality Consistency), and KC (Knowledge Consistency), along with the overall average score. Results are presented for both English and Chinese test sets. End-to-end models are indicated with a gray background. Proprietary models include Alibaba Cloud API and GPT-4o Audio, while all others are open-source.
@@ 4 SpeechRole-Eval
To systematically assess SRPAs, we introduce SpeechRole-Eval, a benchmark designed to evaluate model performance across three key dimensions: interaction ability, speech expressiveness, and role-playing fidelity. Existing evaluations of speech-based role-playing remain largely reliant on human judgments [DBLP:journals/corr/abs-2505-20277], lacking standardized, objective metrics. At the same time, recent research has demonstrated the growing potential of LLMs as automated judges in various evaluation tasks [li2025generation]. Motivated by these developments, SpeechRole-Eval leverages a LLM with speech understanding capabilities to provide automated, objective, and reliable assessment of SRPAs, with the scoring process validated against human annotations. This section describes the task setup, evaluation metrics, and the overall evaluation pipeline.